two small ball having the same mass and charge and located at on same vertical at height H1 and H2 shown in the same direction along a horizontal at same velocity v and first ball thrown the ground at a distance l from initial what is his at what height will the second wall be at the instant the air dragand the charges and induced should be neglected

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
 

let mass of both the charges is m and charge on both the charges is q

take a sphere which is at height h1 two forces act on it

force of gravitation in downword direction and electrostatic force in downward direction hence net downward direction is qe+mg/m hence accelaration is qe/m+g hence

h1=1/2(g+qe/m)t^2

let t is time taken by small sphere in travelling distance

l=vt , t=l/v

let x is distance travelled by second sphere

forces acting on it

force of gravitation downward and electrostatic force upwards hence net acceleration in downward direction is( g-qe/m)

X=1/2(g-qe/m)t^2

height at this instant is h2-x

x=gl^2/v^2-h1

hence height is h2-(gl^2/v^2)

 

 

 

Regards

Arun (askIITians forum expert)